A 500 kg roller coaster is pulled to the top of a hill 45 m high and arrives at the top with a zero speed.



a) how much work has done by the vhain to pull the car to the top?



b) how much potential energy does it have at the top?



c)What is the total energy of the system?



d) how much kinetic energy does it have at the bottom?



1
Expert's answer
2022-02-22T08:52:50-0500

Given:

"m=500\:\rm kg"

"h=45\:\rm m"

"g=9.8\:\rm m/s^2"


a) how much work has done by the vhain to pull the car to the top?


"W=Fd\cos\theta=mgh=500*9.8*45=220500\:\rm J"

b) how much potential energy does it have at the top?


"E_p=mgh=500*9.8*45=220500\:\rm J"

c)What is the total energy of the system?

"E=E_k+E_p=\frac{mv^2}{2}+mgh=0+220500=220500\:\rm J"

d) how much kinetic energy does it have at the bottom?

"E_k=E-E_p=220500-0=220500\:\rm J"